# The impact of hyperons on neutron star mergers: gravitational waves, mass ejection and black hole formation

### v1

@@ II Thermal behavior of the hyperonic Equation of State
As a subgroup of the covariant models, the SU(3) Chiral Mean Field (CMF) is also considered, on which the DNS model ([102]) is based. Using the spontaneous chiral symmetry breaking, the masses of the baryon octet are generated via the coupling to the scalar meson field. At high temperatures/densities the chiral symmetry is restored, and the effective masses of the baryons are significantly lowered.
@@
As we motivated in the introduction, it is important to investigate the thermal contribution to the pressure and the energy density. The thermal contribution to a quantity $X$ is defined as the difference between the quantity at some finite temperature $T$, density $\rho_{B}$ and charge fraction $Y_{Q}$ and the quantity at the same density and charge fraction but at zero temperature: $X_{th}=X(T,\rho_{B},Y_{Q})-X(0,\rho_{B},Y_{Q})$.
Figure 2: Thermal energy density $\epsilon_{\mathrm{th}}$ and thermal pressure $P_{\mathrm{th}}$ as a function of density at fixed charge fraction $Y_{Q}=0.1$ and temperature $T=25$ MeV. For comparison, with black curves, three different nucleonic models (solid line for SFHO, dashed line for FSU2R, dash-dotted line for DD2) are also shown.
@@ II.1 Density dependence of thermodynamical quantities at finite temperature
Figure 3: $\Gamma_{\mathrm{th}}$ (left plot) and $\Delta\rho_{Y}$ (right plot) as a function of density at fixed charge fraction $Y_{Q}=0.1$ and temperature $T=25$ MeV. In the left panel, for comparison with black curves, three different purely nucleonic models (solid line for SFHO, dashed line for FSU2R, dash-dotted line for DD2) are also shown.
@@ III.2.1 Relative frequency shifts
To assess the influence of the finite-temperature regime of the EoS, we focus our analysis on the post-merger phase and its GW emission. In our previous study [89] we discussed that for the same cold EoS the presence of hyperons at finite temperature yields an increase of the dominant post-merger GW frequency compared to the case when only nucleons are considered. In order to quantify this shift, we define $\Delta f=f_{\mathrm{peak}}-f_{\mathrm{peak}}^{1.75}$ with $f_{\mathrm{peak}}$ being the dominant GW frequency from the simulation with the fully temperature dependent EoS table and $f_{\mathrm{peak}}^{1.75}$ being the dominant frequency from the $\Gamma_{\mathrm{th}}$-run, i.e. the calculation with the zero-temperature beta-equilibrium EoS slice of the same model and an approximate thermal treatment using $\Gamma_{\mathrm{th}}=1.75$. As argued above, the choice of $\Gamma_{\mathrm{th}}=1.75$ is intentional since it has been shown to reproduce the thermal behavior of nucleonic models, which is fully corroborated by our analysis comparing full tables and $\Gamma_{\mathrm{th}}$-runs for purely nucleonic models (see below and [89]). In this sense, the quantity $\Delta f$ is a measure of how much a given model deviates from an “idealized” nucleonic thermal behavior.
@@ III.2.2 $f_{\mathrm{peak}}-\Lambda$ relations
We now consider $f_{\mathrm{peak}}-\Lambda_{\mathrm{M}}$ relations for the set of EoS models considered here, which is larger than the sample in [89]. Following the approach of earlier works [89, 131], we first identify the stellar mass $M$ for which the $f_{\mathrm{peak}}-\Lambda_{M}$ relation of purely nucleonic models becomes particularly tight. In Fig. 7 we show the maximum and the average deviation of the data from a quadratic fit $f_{\mathrm{peak}}(\Lambda_{M})$, for purely nucleonic EoSs, as functions of the mass $M$. The average scatter with respect to a quadratic fit is almost constant for different reference masses $M$. The maximum deviation shows a clear minimum at $M=1.65\leavevmode\nobreak\ M_{\odot}$, which we thus choose as a reference value [^5] Note that in [89] the reference mass was 1.75 $M_{\odot}$ for equal-mass binaries, whereas the larger set of EoS models considered here suggests a slightly lower reference mass..
@@ III.4 Further analysis
Figure 18: Average temperature of the remnant as a function of time for all hyperonic models. For comparison, three nucleonic models (DD2 - dotted line, SFHO - solid line, FSU2R - dashed line) are also shown with black lines. The vertical lines show the characteristic time window that we use to compute $\bar{\Gamma}_{\mathrm{th}}$.
Figure 19: Mass and time averaged temperature inside of the remnant 10 ms after merging as a function of the radius of a $1.4M_{\odot}$ NS.
@@
Finally, we quantitatively assess the relation between EoS stiffness and the average temperature in Fig. 19. We extract the mass and time averaged temperature in the remnant 10 ms after merging in a time window of 2.5 ms. We plot the average temperature as function of the radius of a $1.4M_{\odot}$ star as proxy for the EoS stiffness. Fig. 19 illustrates that the average temperature roughly anticorrelates with the NS radius and that hyperonic models yield temperatures which are reduced compared to purely nucleonic models with the same radius, i.e. EoS stiffness.
Figure 20: Ejecta mass $M_{\mathrm{ej}}$ as a function of the radius $R_{1.4}$ of 1.4 $M_{\odot}$ NS. Crosses display results obtained for models with hyperons and asterisks mark models that additionally include $\Delta$ resonances. The coloring indicates whether the heavy baryons represent more than 1% of the baryon population (green points) or not (blue points). Nucleonic models are labeled with pluses.
@@ III.5 Mass ejection
In Fig. 20 we show the ejecta mass as function of the radius $R_{1.4}$ of a 1.4 $M_{\odot}$ NS as an approximate measure for the EoS stiffness. We employ the same notation as in previous plots with the coloring indicating the number fraction of heavy baryons (green data points denoting models with more than one per cent; blue symbols for models with less than one per cent; black plus signs for purely nucleonic EoSs). There is a clear trend for softer EoSs, i.e. EoSs with smaller radii, to produce significantly higher ejecta masses (cf. [158, 159]). Most hyperonic models follow the behavior of purely nucleonic simulations. Some of the hyperonic models, however, yield ejecta masses which are up to a factor two increased compared to nucleonic models with the same $R_{1.4}$. Such a strong increase represents a very significant effect and may be used to discriminate nucleonic and hyperonic models in future measurements. This requires to better understand mass ejection and kilonova emission in general, which are still subject to various uncertainties, and to determine the dependence of the ejecta increase on the hyperon content.
@@ V Simulations of asymmetric binaries
The results from the simulations are given in Table 5, while Fig. 22 shows the relative frequency shift $\Delta f/f_{\mathrm{peak}}$ as a function of both the peak frequency $f_{\mathrm{peak}}$ (left panel) and the average thermal index of the remnant $\Gamma_{th,av}$ (right panel). The frequency shift, which is characteristic of EoSs with exotic baryons, is clearly visible also for asymmetric binaries. For some of the EoSs with hyperons and $\Delta$ baryons, the relative shift in asymmetric mergers is even more enhanced, reaching values up to 6% as compared to at most 4% for equal-mass systems. Note that only 13 different EoS data points are displayed in Fig. 22, although we consider 14 EoS models with heavy baryons in our calculations. The reason is that, for the softest EoS model (SFHOY), the remnant undergoes a prompt collapse in the simulation with the full 3D EoS table. In the corresponding simulation with $\Gamma_{\mathrm{th}}=1.75$ the central object does not immediately form a black hole. This different qualitative behavior is another clear signature of the effects that thermal hyperons have on the remnant. The influence of the thermal behavior of EoSs with heavy baryons on the threshold mass will be discussed in Section VI. In addition, some of the EoSs produce two dominant peaks with a comparable strength (those points are connected with a green dashed lines). The exact mechanism that produces these peaks is still unknown.
@@ VI Threshold mass for prompt collapse
The promising aspect of this intricacy is, however, the prospect that $M_{\mathrm{thres}}$ may reveal the presence of heavy baryons if measured well enough. In Fig. 28 we compare the threshold mass resulting from the simulations with the fully temperature dependent EoS tables and with the approximate thermal treatment. Hyperons can reduce $M_{\mathrm{thres}}$ by about 0.05 $M_{\odot}$ in comparison to the threshold mass one would expect for a purely nucleonic system with the same stellar parameters of cold NSs. [^9] One may also conclude from Fig. 28 that choosing $\Gamma_{\mathrm{1.75}}$ for purely nucleonic EoSs to determine $M_{\mathrm{thres}}$ may yield a slight underestimation of the threshold mass. Note that inferring the occurrence of hyperons would require a well measured $M_{\mathrm{thres}}$ as well as precise knowledge of $M_{\mathrm{max}}$ and $R_{1.6}$ to compare the measured threshold mass with the one expected for nucleonic EoSs.

### v2

@@ II Thermal behavior of the hyperonic Equation of State
As a subgroup of the covariant models, the SU(3) Chiral Mean Field (CMF) is also considered, on which the DNS model [102] is based. Using the spontaneous chiral symmetry breaking, the masses of the baryon octet are generated via the coupling to the scalar meson field. At high temperatures/densities the chiral symmetry is restored, and the effective masses of the baryons are significantly lowered.
@@
As we motivated in the introduction, it is important to investigate the thermal contribution to the pressure and the energy density. The thermal contribution to a quantity $X$ is defined as the difference between the quantity at some finite temperature $T$, density $\rho_{B}$ and charge fraction $Y_{Q}$ and the quantity at the same density and charge fraction but at zero temperature: $X_{\mathrm{th}}=X(T,\rho_{B},Y_{Q})-X(0,\rho_{B},Y_{Q})$.
Figure 2: Thermal energy density $\epsilon_{\mathrm{th}}$ and thermal pressure $P_{\mathrm{th}}$ as functions of density at fixed charge fraction $Y_{Q}=0.1$ and temperature $T=25$ MeV. For comparison, with black curves, three different nucleonic models (solid line for SFHO, dashed line for FSU2R, dash-dotted line for DD2) are also shown.
@@ II.1 Density dependence of thermodynamical quantities at finite temperature
Figure 3: $\Gamma_{\mathrm{th}}$ (left plot) and $\Delta\rho_{Y}$ (right plot) as functions of density at fixed charge fraction $Y_{Q}=0.1$ and temperature $T=25$ MeV. In the left panel, for comparison with black curves, three different purely nucleonic models (solid line for SFHO, dashed line for FSU2R, dash-dotted line for DD2) are also shown.
@@ III.2.1 Relative frequency shifts
To assess the influence of the finite-temperature regime of the EoS, we focus our analysis on the post-merger phase and its GW emission. In our previous study [89] we discussed that for the same cold EoS the presence of hyperons at finite temperature yields an increase of the dominant post-merger GW frequency [^5] In this study we determine GW frequencies by computing $h_{\mathrm{eff,\times}}(f)=f\cdot\tilde{h}_{\times}(f)$ with the Fourier transform $\tilde{h}_{\times}(f)$ of the strain of the cross polarization. Since the spectrum $h_{\mathrm{eff,\times}}(f)$ is given by discrete data points, we determine maxima of peaks by fitting a parabolic function to the three highest data points at the top of a peak. From the analytic representation we obtain the location of the maximum of the respective peak. compared to the case when only nucleons are considered. In order to quantify this shift, we define $\Delta f=f_{\mathrm{peak}}-f_{\mathrm{peak}}^{1.75}$ with $f_{\mathrm{peak}}$ being the dominant GW frequency from the simulation with the fully temperature dependent EoS table and $f_{\mathrm{peak}}^{1.75}$ being the dominant frequency from the $\Gamma_{\mathrm{th}}$-run, i.e. the calculation with the zero-temperature beta-equilibrium EoS slice of the same model and an approximate thermal treatment using $\Gamma_{\mathrm{th}}=1.75$. As argued above, the choice of $\Gamma_{\mathrm{th}}=1.75$ is intentional since it has been shown to reproduce the thermal behavior of nucleonic models, which is fully corroborated by our analysis comparing full tables and $\Gamma_{\mathrm{th}}$-runs for purely nucleonic models (see below and [89]). In this sense, the quantity $\Delta f$ is a measure of how much a given model deviates from an “idealized” nucleonic thermal behavior.
@@ III.2.2 $f_{\mathrm{peak}}-\Lambda$ relations
We now consider $f_{\mathrm{peak}}-\Lambda_{\mathrm{M}}$ relations for the set of EoS models considered here, which is larger than the sample in [89]. Following the approach of earlier works [89, 131], we first identify the stellar mass $M$ for which the $f_{\mathrm{peak}}-\Lambda_{M}$ relation of purely nucleonic models becomes particularly tight. In Fig. 7 we show the maximum and the average deviation of the data from a quadratic fit $f_{\mathrm{peak}}(\Lambda_{M})$, for purely nucleonic EoSs, as functions of the mass $M$ [^6] Considering all purely nucleonic EoSs, we perform a quadratic least-squares fit $f_{\mathrm{peak}}(\Lambda_{M})$ and compute the residual between the fit and the true value as $r_{i}=|f_{\mathrm{peak}}-f_{\mathrm{fit}}|$ for each data point. For every fixed reference mass $M$ we compute the average value of the residuals $r_{\mathrm{ave}}=\sum r_{i}/N$ with $N$ being the number of data points and determine the maximum of all residuals $r_{\mathrm{max}}=\max{(r_{i})}$.. The average scatter with respect to a quadratic fit is almost constant for different reference masses $M$. The maximum deviation shows a clear minimum at $M=1.65\leavevmode\nobreak\ M_{\odot}$, which we thus choose as a reference value [^7] Note that in [89] the reference mass was 1.75 $M_{\odot}$ for equal-mass binaries, whereas the larger set of EoS models considered here suggests a slightly lower reference mass..
@@ III.4 Further analysis
Figure 18: Average temperature of the remnant as a function of time for all hyperonic models. For comparison, three nucleonic models (DD2 - dotted line, SFHO - solid line, FSU2R - dashed line) are also shown with black lines. The vertical lines show the characteristic time window that we use to compute $\bar{\Gamma}_{\mathrm{th}}$ and $\bar{T}$
Figure 19: Mass and time averaged temperature inside of the remnant 5 ms after merging as a function of the radius of a $1.4M_{\odot}$ NS.
@@
Finally, we quantitatively assess the relation between EoS stiffness and the average temperature in Fig. 19. We extract the mass and time averaged temperature in the remnant 5 ms after merging in a time window of 2.5 ms. We plot the average temperature as function of the radius of a $1.4M_{\odot}$ star as proxy for the EoS stiffness. Fig. 19 illustrates that the average temperature roughly anticorrelates with the NS radius and that hyperonic models yield temperatures which are reduced compared to purely nucleonic models with the same radius, i.e. EoS stiffness.
@@ III.5 Mass ejection
Figure 20: Ejecta mass $M_{\mathrm{ej}}$ as a function of the radius $R_{1.4}$ of 1.4 $M_{\odot}$ NS (left panel). Crosses display results obtained for models with hyperons and asterisks mark models that additionally include $\Delta$ resonances. The coloring indicates whether the heavy baryons represent more than 1% of the baryon population (green points) or not (blue points). Nucleonic models are labeled with pluses. The black curve shows a parabolic fit to the purely nucleonic models. Right panel displays the relative difference between $M_{\mathrm{ej}}$ and a parabolic fit $M_{\mathrm{ej}}^{\mathrm{fit}}(R_{1.4})$ to the purely nucleonic models.
In Fig. 20 (left panel) we show the ejecta mass as function of the radius $R_{1.4}$ of a 1.4 $M_{\odot}$ NS as an approximate measure for the EoS stiffness. We employ the same notation as in previous plots with the coloring indicating the number fraction of heavy baryons (green data points denoting models with more than one per cent; blue symbols for models with less than one per cent; black plus signs for purely nucleonic EoSs). There is a clear trend for softer EoSs, i.e. EoSs with smaller radii, to produce significantly higher ejecta masses (cf. [158, 159]). Most hyperonic models follow the behavior of purely nucleonic simulations. Some of the hyperonic models, however, yield ejecta masses which are up to a factor two increased compared to nucleonic models with the same $R_{1.4}$ (see right panel in Fig. 20). Such a strong increase represents a very significant effect and may be used to discriminate nucleonic and hyperonic models in future measurements. This requires to better understand mass ejection and kilonova emission in general, which are still subject to various uncertainties, and to determine the dependence of the ejecta increase on the hyperon content. The most pronounced increase is found for the simulation with the SFHOY EoS. The likely reason for this strong increase is the fact that the total binary mass for this simulation is very close to the threshold binary for direct black-hole formation (cf. Tab. 5).
@@ V Simulations of asymmetric binaries
The results from the simulations are given in Table 5, while Fig. 22 shows the relative frequency shift $\Delta f/f_{\mathrm{peak}}$ as a function of both the peak frequency $f_{\mathrm{peak}}$ (left panel) and the average thermal index of the remnant $\bar{\Gamma}_{\mathrm{th}}$ (right panel). The frequency shift, which is characteristic of EoSs with exotic baryons, is clearly visible also for asymmetric binaries. For some of the EoSs with hyperons and $\Delta$ baryons, the relative shift in asymmetric mergers is even more enhanced, reaching values up to 6% as compared to at most 4% for equal-mass systems. Note that only 13 different EoS data points are displayed in Fig. 22, although we consider 14 EoS models with heavy baryons in our calculations. The reason is that, for the softest EoS model (SFHOY), the remnant undergoes a prompt collapse in the simulation with the full 3D EoS table. In the corresponding simulation with $\Gamma_{\mathrm{th}}=1.75$ the central object does not immediately form a black hole. This different qualitative behavior is another clear signature of the effects that thermal hyperons have on the remnant. The influence of the thermal behavior of EoSs with heavy baryons on the threshold mass will be discussed in Section VI. In addition, some of the EoSs (hyperonic models from [75]) produce two near-by dominant peaks with a comparable strength (those points are connected with a green dashed line; see Tab. 5). The exact mechanism that produces these peaks is still unknown.
@@ VI Threshold mass for prompt collapse
The promising aspect of this intricacy is, however, the prospect that $M_{\mathrm{thres}}$ may reveal the presence of heavy baryons if measured well enough. In Fig. 28 we compare the threshold mass resulting from the simulations with the fully temperature dependent EoS tables and with the approximate thermal treatment. Hyperons can reduce $M_{\mathrm{thres}}$ by about 0.05 $M_{\odot}$ in comparison to the threshold mass one would expect for a purely nucleonic system with the same stellar parameters of cold NSs. [^11] One may also conclude from Fig. 28 that choosing $\Gamma_{\mathrm{1.75}}$ for purely nucleonic EoSs to determine $M_{\mathrm{thres}}$ may yield a slight underestimation of the threshold mass. Note that inferring the occurrence of hyperons would require a well measured $M_{\mathrm{thres}}$ as well as precise knowledge of $M_{\mathrm{max}}$ and $R_{1.6}$ to compare the measured threshold mass with the one expected for nucleonic EoSs. Also, the exact numerical determination of $M_{\mathrm{thres}}$ depends on its precise definition as well as the physical model, numerical scheme and resolution [168, 167].